OpenStudy (anonymous):

Find the average rate of change of f(x)= 7x^2+30

OpenStudy (anonymous):

@agent0smith @amistre64

OpenStudy (anonymous):

please anyone just need a step by step

OpenStudy (clalgee):

hmm... ok, so do you know where to start off?

OpenStudy (anonymous):

no can you help me @Clalgee can you do it step by step so i can take notes and do my other simmilair question?

OpenStudy (anonymous):

@nincompoop can you help me?

OpenStudy (anonymous):

The Average Rate of Change function describes the average rate at which one quanity is changing with respect to something else changing.

OpenStudy (anonymous):

can you show me step by step so i can do the other one by myself?

OpenStudy (anonymous):

In general, an average rate a change function is a process that calculates the the amount of change in one item divided by the corresponding amount of change in another. Using function notation, we can define the Average rate of Change of a function f from a to x as \[A(x)=\frac{ f(x)-f(a) }{ x-a }\]

OpenStudy (anonymous):

A is the name of this average rate of change function x - a represents the change in the input of the function f f(x) - f(a) represents the change in the function f as the input changes from a to x a few examples to start: Find the slope of the line going through the curve as x changes from 3 to 0. Step 1: f (3) = -1 and f (0) = -4 Step 2: Use the slope formula to create the ratio \[\frac{ f(0)-f(3) }{ 0-3 }\] Step 3: Simplify. \[\frac{ f(0)-f(3) }{ 0-3 }=\frac{ -4-(-1) }{ 0-3 }=1\] Step 4: So the slope of the line going through the curve as x changes from 3 to 0 is 1.

OpenStudy (anonymous):

ohh okay so would the rate of change be increasing or decreasing?

OpenStudy (anonymous):

sec...

OpenStudy (anonymous):

fo can you tell me where my 7x^2 and 30 fit into the formula?

OpenStudy (anonymous):

had to get my graphing calc

OpenStudy (anonymous):

in your question do they state x=? and x=?

OpenStudy (anonymous):

find the AROC of (7x^2)+30 from x= something to x= something

OpenStudy (anonymous):

let me show you my question i have graph also

OpenStudy (anonymous):

that is a weird question for sure

OpenStudy (anonymous):

yes i dont understadn it i have the graph i need to find the rate of change and see if its increasing or decreasing

OpenStudy (anonymous):

let me find my old book and see what it says brb

OpenStudy (anonymous):

so i think they want you to use the above formula, and substitute x values

OpenStudy (anonymous):

for instance when x= -0.5 and x=1

OpenStudy (anonymous):

that will give you an AROC

OpenStudy (anonymous):

\[AROC= \frac{ f(-0.5)-f(1) }{ (-0.5)-1() }\]

OpenStudy (anonymous):

f(-0.5)= 7((-0.5)^2)+30 = and f(1)= 7((1)^2)+30 =

OpenStudy (anonymous):

\[\frac{ f(-0.5)-f(1) }{ -0.5-1 }=\frac{31.75-37 }{ -0.5-1 }=\frac{ -5.25 }{-1.5 }\]

OpenStudy (anonymous):

AROC = 3.5

OpenStudy (anonymous):

increasing with justification :)

OpenStudy (anonymous):

you can try it again with any other x values, for practice, then substitute into your equation

OpenStudy (anonymous):

oh k tahnks :)

OpenStudy (anonymous):

np